# HADIS: Hybrid Adaptive Diffusion Model Serving for Efficient Text-to-Image Generation

### v1

@@ 4.2. Performance on Real-World Traces
4.3. Performance on Synthetic Traces
@@
In Figure 6(b)-left, we can observe per-model throughputs (processed queries per second, QPS) and percentages of queries shared by each model. The traffic mix shifts toward faster models (i.e., SDXL-Lightning and SD3.5-Turbo) during P3/P4 and back toward higher-quality models (i.e., SD3.5-Medium and SD3.5-Large¿) during P1. This is because the MILP chooses model combinations from the profiled table that meet the active SLO under the current load. When QPS rises, HADIS responses by adjusting the hardness and quality threshold to reduce the fraction of query re-routing, leading to an increased percentage of queries on lightweight paths. The observation is more obvious in Figure 6(b)-right, which shows the GPU workers allocated and the ratio of queries routed to the heavyweight model. Note that in P3 and P4, HADIS is able to achieve high quality with heavier models while meeting latency SLO, because SDXL-Lightning is applied periodically to drain the queued queries accumulated by those models, preventing timeouts.
@@
Figure 7. Effectiveness of hybrid routing. The hybrid routing (HADIS) outperforms Random and Router-only (Router) in FID and reduces SLO violations compared to Discriminator-only (Disc), achieving the best quality-latency trade-off under identical resources.
4.4. Effectiveness of Hybrid Routing
@@
We now show an ablation study of our hybrid routing approach in Figure 7 to understand the effect of the rule-based router and the image discriminator. We isolate the routing policy by running four variants: (1) Random: we randomly assign hardness and quality scores to each query, using these random scores for routing decision making. (2) Router-only: the routing decisions are made by the rule-based router only without discriminator-based re-routing. (3) Discriminator-only: the routing decisions are made by the discriminator only without the router. (4) HADIS: hybrid routing is enabled with both rule-based query routing and discriminator-based re-routing. All variants are evaluated on the same real-world trace with the same MILP-based resource allocator.
@@
Figure 7 shows that HADIS achieves the lowest average FID and the lowest SLO violation ratio across all variants. Random and Router-only suffer from worse FID at similar SLO violation levels since hard queries are served by lightweight models due to misclassification. Discriminator-only improves FID but suffers from higher violations, as it forces every query through the light model, paying extra latency. In contrast, HADIS adopts hybrid routing, combining query-aware rules to skip the light stage for obviously hard queries and a learned discriminator to catch false “easy” cases, while adjusting thresholds to match workload. Thus, it preserves low latency and achieves higher response quality.
@@
Figure 8. Comparison of resource allocation strategies. HADIS applies MILP. Cache-{D} caches plans using demands as key; Cache-{D+Q} caches plans using demands and queue length as key. Cached policies lag during load shifts, showing transient quality drops and violation spikes, whereas HADIS adapts in real time, yielding lower violations with comparable or better FID.
4.5. Evaluation of Resource Allocation
@@
We ablate the resource planner by implementing two variants, Cache-{D+Q} and Cache-{D}. Unlike HADIS which solves an online MILP to seek optimal resource allocation plans at runtime, these cached planners select and apply a previously computed resource allocation configuration on a cache hit. Cache-{D} stores plans in a lookup table keyed by a discretized range of system demand, whereas Cache-{D+Q} uses a two-dimensional key composed of a range of system demand and a range of worker queue length. At runtime, we quantize the observed demand (and queue length for Cache-{D+Q}) to their respective ranges and probe the table. If the measurements fall within any keyed range that has an associated plan, we declare a cache hit and apply that plan. If not, we solve the MILP to obtain a plan and then insert it into the lookup table under the corresponding range key.
@@
Figure 8 shows the timeseries results and average statistics. Cache-{D} performs the worst, where high FID and SLO violation spikes early in the run were observed. The planner lags during load shifts as it is agnostic about the real-time system workload. Cache-{D+Q} is more responsive than Cache-{D} by capturing congestion through the extra key dimension in the lookup table. We empirically observe that both caching variants exhibit a startup spike in SLO violations since their selected plans are either biased or unsynchronized with the actual real-time demands. Once the workload settles, violations stabilize for both variants.
Unlike the cached plans that react only when the key changes, HADIS with online MILP adapts plans in real time, rebalancing workers and thresholds as demand changes. It prevents prolonged queue growth and keeps low violations while maintaining comparable or better FID, avoiding under- or over-provision when workload fluctuates.
@@
Figure 9. Sensitivity to latency SLO on HADIS and baselines. HADIS guarantees low SLO violations and high quality over varying SLO values across all methods.
4.6. Sensitivity to Latency SLO
@@
4.7. Sensitivity to Hardness of Queries
@@
We study how sensitive HADIS is to shifts in the hardness distribution of incoming queries. We isolate the effect of hardness from demand by fixing the query demand and constructing a four-phase trace: the first and third phases contain only Easy queries, and the second and fourth contain only Hard queries. Queries are categorized using weighted scores over textual features. As shown in Figure 10, during Easy phases, the system settles into a relatively dynamic phase where the lightweight model serves the majority of traffic. Immediately after each transition to a Hard phase, the router bypasses lightweight inference for a larger fraction of queries, sharply increasing the heavy-class intake. The heavy queue rises, triggering workers assigned to the heavyweight model. These results confirm that HADIS adaptively concentrates resources on heavyweight models precisely when hard queries predominate, highlighting the value of our query-side bypass design.
@@
Figure 10. Sensitivity to Query Hardness. HADIS adaptively bypasses lightweight inference and reallocates workers according to query complexity.
4.8. Overhead Analysis
@@
HADIS introduces three sources of overhead: (1) a query router that performs rule-based feature extraction and a score computation (5ms per query), (2) a CLIP-based discriminator that computes quality score for generated images (7ms per query on an L40S GPU), and (3) an MILP-based resource allocator whose measured average runtime is $\sim$30ms per planning cycle. The aggregated overhead of router and discriminator is $\sim$12ms, which is negligible relative to diffusion inference of SDXL-Lightning (2.4%), SD3.5-Turbo (0.9%), SD3.5-Medium (0.01%), and SD3.5-Large (¡0.01%). Moreover, since the MILP solver runs periodically and applies updates asynchronously, this overhead is off the critical path and does not affect individual queries.
5. Related Work
@@
Model serving systems aim to provide a unified abstraction to the user, hiding details of the underlying ML frameworks, data preprocessing, and performance optimization. Representative production systems include SageMaker (SageMaker, 2020), Triton (Tri, 2022), TensorFlow Serving (Ten, 2024), and TorchServe (PyT, 2023), whereas research prototypes include Nexus (Shen et al., 2019), BATCH (Ali et al., 2020), and Clockwork (Gujarati et al., 2020). While mature and widely deployed, these systems typically require users to specify the models and then handle resource management based on the user-specified models. Without reasoning about multiple model variants for the same task, such framework gives up opportunities to further balance accuracy, latency, and cost. Some recent systems improve upon it by managing both resources and model variants. For instance, Clipper (Crankshaw et al., 2017b), Rafiki (Wang et al., 2018), and Cocktail (Gunasekaran et al., 2022) use model ensembles to enhance response quality. INFaaS (Romero et al., 2021), Model Switching (Zhang et al., 2020), and Sommelier (Guo et al., 2022) adaptively select model variants in response to system loads. Proteus (Ahmad et al., 2024) exploits accuracy–efficiency trade-offs across model variants to meet throughput requirements while maximizing accuracy. Still, systems that primarily react to system state and ignore query difficulty miss optimization opportunities when requests vary in semantic complexity. A cascading architecture for large language model (LLM) serving has therefore been proposed, such as CascadeBERT (Li et al., 2020) and Tabi (Wang et al., 2023). These systems allow easy queries to be processed exclusively by light models and decide whether to return results or re-route inputs to more complex models. DiffServe (Ahmad et al., 2025) constructs diffusion model cascades and frames resource allocation as a mixed integer linear programming (MILP) problem to optimize both the efficiency and response quality. Still, the cascading frameworks that fixes the cascade pair and always initiates with a lightweight model pass can wast compute of re-routing difficult prompts. HADIS improves upon above frameworks by adaptively selecting the optimal model pair and decision thresholds based on the time-varying query demands. Unnecessary lightweight inference is also avoided by sending obviously hard prompts directly to heavyweight models.
6. Conclusion
We presented HADIS, a diffusion model serving system that couples adaptive model cascading with hybrid routing and a lookup-table-driven MILP planner to meet latency SLOs while delivering high response quality. HADIS profiles candidate model pairs offline to build Pareto-optimal cascade configurations, then selects the best pair and thresholds online while jointly allocating GPUs and batch sizes. A rule-based router and a CLIP-based discriminator work in tandem to preemptively route clearly hard queries and validate outputs from lightweight models, reducing cascade overhead without sacrificing quality. Evaluated on real-world workload traces, HADIS outperforms state-of-the-art baselines, improving response quality and reducing SLO violations.

### v2

@@ 5.2 Performance on Real-World Traces
5.3 Performance on Synthetic Traces
@@
In Figure 8(b)-left, we can observe per-model throughputs (processed queries per second, QPS) and percentages of queries shared by each model. The traffic mix shifts toward faster models (i.e., SDXL-Lightning and SD3.5-Turbo) during P3/P4 and back toward higher-quality models (i.e., SD3.5-Medium and SD3.5-Large>) during P1. This is because the MILP chooses model combinations from the profiled table that meet the active SLO under the current load. When QPS rises, HADIS responses by adjusting the hardness and quality threshold to reduce the fraction of query re-routing, leading to an increased percentage of queries on lightweight paths. The observation is more obvious in Figure 8(b)-right, which shows the GPU workers allocated and the ratio of queries routed to the heavyweight model. Note that in P3 and P4, HADIS is able to achieve high quality with heavier models while meeting latency SLO, because SDXL-Lightning is applied periodically to drain the queued queries accumulated by those models, preventing timeouts.
@@
Figure 9: Effectiveness of hybrid routing. The hybrid routing (HADIS) outperforms Random and Router-only (RO) in FID and reduces SLO violations compared to Discriminator-only (DO), achieving the best quality-latency trade-off under identical resources.
5.4 Effectiveness of Hybrid Routing
@@
We now show an ablation study of our hybrid routing approach in Figure 9 to understand the effect of the rule-based router and the image discriminator. We compare HADIS against three variants that disable parts of the hybrid routing pipeline: (1) Random: assign random hardness and quality scores to each query and routes based on these scores. (2) Router-only: routing decisions are made by the rule-based router only without discriminator-based re-routing. (3) Discriminator-only: routing decisions are made by the discriminator only without the router. All variants are evaluated on the same real-world trace with the same MILP-based resource allocator.
@@
Figure 9 shows that HADIS achieves the lowest average FID and SLO violation ratio across all variants. Random and Router-only suffer worse FID at similar SLO violation levels since many hard queries are served by lightweight models due to misclassification. Discriminator-only improves FID but incurs higher violations, as it forces every query through the light model, paying extra latency. In contrast, HADIS adopts hybrid routing, skipping the light stage for obviously hard queries and catching false “easy” cases with discriminator. Thus, it preserves low latency and achieves higher response quality.
@@
Figure 10: Comparison of resource allocation strategies. HADIS applies MILP. Cache-{D} caches plans using demands as key; Cache-{D+Q} caches plans using demands and queue length as key. Cached policies lag during load shifts, showing transient quality drops and violation spikes, whereas HADIS adapts in real time, yielding lower violations with comparable or better FID.
5.5 Evaluation of Resource Allocation
@@
We ablate the resource planner by implementing two variants, Cache-{D+Q} and Cache-{D}. Unlike HADIS which solves an online MILP to seek optimal resource allocation plans at runtime, these cached planners reuse previously computed plans on a cache hit. Cache-{D} indexes plans by a discretized range of system demand, whereas Cache-{D+Q} uses a two-dimensional key over range of system demand and range of worker queue length. At runtime, we quantize the observed demand (and queue length for Cache-{D+Q}) to their ranges and probe the cache. If the measurements fall within any keyed range that has an associated plan, we declare a cache hit and apply that plan. If not, we solve the MILP and insert the new plan under the corresponding range key.
@@
Figure 10 shows the timeseries results and average statistics. Cache-{D} performs the worst, with early spikes in FID and SLO violations because it lags behind load shifts and is agnostic about real-time system workload. Cache-{D+Q} is more responsive by incorporating queue length into the key to capture congestion. We empirically observe that both caching variants exhibit a startup spike in SLO violations since their initial plans are either biased or misaligned with the actual demands, but violations stabilize once the workload settles. In contrast, HADIS with online MILP adapts plans in real time, rebalancing workers and thresholds as demand changes. It prevents prolonged queue growth, keeps violations low, and maintains comparable or better FID, avoiding under- or over-provision when workload fluctuates.
@@
Figure 11: Sensitivity to latency SLO on HADIS and baselines. HADIS guarantees low SLO violations and high quality over varying SLO values across all methods.
5.6 Sensitivity to Latency SLO
@@
5.7 Sensitivity to Hardness of Queries
@@
We study how sensitive HADIS is to shifts in the hardness distribution of incoming queries. We isolate the effect of hardness from demand by fixing the query demand and constructing a four-phase trace: the first and third phases contain only Easy queries, and the second and fourth contain only Hard queries. Queries are categorized using weighted scores over textual features. As shown in Figure 12, during Easy phases, the system settles into a relatively dynamic phase where the lightweight model serves the majority of traffic. At each transition to a Hard phase, the router bypasses lightweight inference for a larger fraction of queries, sharply increasing the heavy-class intake. The heavy queue rises, triggering workers assigned to the heavyweight model. These results confirm that HADIS adaptively concentrates resources on heavyweight models precisely when hard queries predominate, highlighting the value of our query-side bypass design.
@@
Figure 12: Sensitivity to Query Hardness. HADIS adaptively bypasses lightweight inference and reallocates workers according to query complexity.
5.8 Overhead Analysis
@@
HADIS introduces three sources of overhead: (1) a query router that performs rule-based feature extraction and a score computation (5ms per query), (2) a CLIP-based discriminator that computes quality score for generated images (7ms per query on an L40S GPU), and (3) an MILP-based resource allocator whose measured average runtime is $\sim$30ms per planning cycle. The aggregated overhead of router and discriminator is $\sim$12ms, which is negligible relative to diffusion inference of SDXL-Lightning (2.4%), SD3.5-Turbo (0.9%), SD3.5-Medium (0.01%), and SD3.5-Large (<0.01%). Moreover, since the MILP solver runs periodically and applies updates asynchronously, this overhead is off the critical path and does not affect individual queries.
6 Related Work
@@
Model serving systems aim to provide a unified abstraction to the user, hiding details of the underlying ML frameworks, data preprocessing, and performance optimization. Systems such as SageMaker [36], Triton [1], TensorFlow Serving [3], TorchServe [2], and research prototypes like Nexus [39], BATCH [10], and Clockwork [16] typically require users to manually specify which models to deploy while the system handles resource management accordingly. More recent systems tend to jointly optimize accuracy, latency, and cost together for more efficient resource allocation. Clipper [14], Rafiki [42], and Cocktail [17] ensemble models to improve quality, while INFaaS [34], Model Switching [45], and Sommelier [18] adaptively select variants under changing load. Proteus [7] further exploits accuracy–efficiency trade-offs across variants. However, these systems react primarily to system state rather than per-query difficulty, missing optimization opportunities when requests vary in semantic complexity. Cascaded architectures have therefore been proposed (e.g., CascadeBERT [25], Tabi [43]), where easy queries are handled by lightweight models and harder ones are deferred to heavier models. DiffServe [8] extends this idea to diffusion models, constructing model cascades and formulating resource allocation as an MILP to jointly optimize efficiency and response quality. NIRVANA [6] reduces cost by caching intermediate denoising states, which is complementary to our approach. All these cascading frameworks, however, fix the cascade pair and always begin with a lightweight pass, which wastes computation when many prompts are clearly difficult. In contrast, HADIS adaptively selects both the model pair and decision thresholds based on time-varying query demands, and uses a prompt-side router to send obviously hard prompts directly to heavyweight models, thereby improving generation quality and reducing cost by avoiding unnecessary lightweight inference.
7 Conclusion
In this work, we presented HADIS, a query-aware text-to-image diffusion model serving system that addresses the fundamental tension between response quality, latency, and resource efficiency in diffusion model deployments. It advocates a hybrid routing architecture that integrates a predictive router with a generation-aware discriminator, reducing unnecessary computation on hard queries while preserving high-quality outputs for easy ones. HADIS reduces resource management complexity through offline profiling and a Pareto-optimal lookup table, and jointly optimize architectural and resource decisions with an MILP formulation. Evaluated on real-world workload traces, HADIS outperforms state-of-the-art baselines, improving response quality and reducing SLO violations.